40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Nov 11 22:11:15 2015, Koji, Update, SUS, PSL/IOO maintenance, TM SUS check up TM_SUS.pdfTM_SUS_SD.pdf
    Reply  Thu Nov 12 09:15:56 2015, Steve, Update, SUS, ITMX osems ITMX.png
       Reply  Fri Nov 13 15:20:24 2015, Steve, Update, SUS, OSEMs OSEMs.pngPOSbiasOFF.png
    Reply  Thu Nov 12 09:41:02 2015, Steve, Update, PSL, PMC PMC_ITMX.png
    Reply  Tue Nov 17 00:57:21 2015, ericq, Update, SUS, ITMX UL calmed? TM_nov17.pdfITMX_sensortrend.png
       Reply  Tue Nov 17 15:59:23 2015, rana, Update, SUS, ITMX UL calmed? 
Message ID: 11753     Entry time: Wed Nov 11 22:11:15 2015     Reply to this: 11755   11756   11770
Author: Koji 
Type: Update 
Category: SUS 
Subject: PSL/IOO maintenance, TM SUS check up 

PMC

- Before any of the following work, I went to the PSL table and aligned the PMC. In fact, it has not been misaligned.


IMC

- It was claimed in the meeting today that the IMC had not been happy thesedays. I checked out what's happening.

- I found the IMC was still well aligned. Autolocker frequently stuck on a weak higher-order mode and couldn't recover TEM00 locking without help.

- I modified /opt/rtcds/caltech/c1/scripts/MC/mcdown for easier relocking on TEM00.

The MC_REFL_GAIN and MC_VCO_GAIN for relocking was set to be 27 and -3, in stead of 0 and 10, respectively.
This means that REFL_GAIN is not changed before and after the locking. Only VCO_GAIN is lowered for lock acquisition.
The corresponding lines in mcdown are excerpted here.

#set servo and boost gains for re-acquisition
#${ewrite} C1:IOO-MC_REFL_GAIN 0 &
${ewrite} C1:IOO-MC_REFL_GAIN 27 &
#${ewrite} C1:IOO-MC_VCO_GAIN 10 &
${ewrite} C1:IOO-MC_VCO_GAIN -3 &

We still have some chance of locking on higher-order modes. If I jiggle the VCO gain slider from -31 to 0, eventually I find TEM00.
I don't know how to do it in the script yet. For now, I increased tickle amplitude from 300 to 500.

/opt/rtcds/caltech/c1/scripts/MC/MC2tickleON

amp=300
=>

amp=500

- IMC was locked and aligned with the WFS. The WFS feedback offsets were offloaded to the alignment slider (via the MEDM button as usual).


TM SUS

- I wanted to use ASS. => The OL damping and ASC inputs are enabled for ETMX. The filter bank output of the ASS servos are all turned on.
- The arms were locked and aligned with ASS. The ASS servo offsets were offloaded to the ASC offset sliders (as usual).

- I found the X arm spot positionis moving slowly in pitch. I wanted to know what is causing this.

- Turned off all the OPLEV dampings for the four test masses.
- Took the power spectra of the OSEM output (e.g. C1:SUS-ITMX_LLSEN_OUT)

See attachment 1 (The DTT XML file can be found as /users/koji/151111/TM_SUS.xml )

- It seems that something is wrong with ITMX UL OSEM
  The signal level seems to be identical with the others. However, the noise level is huge. We need to check if the cable connection is OK.

- ITMY LL shows remarkably higher bounce mode although I can't tell if this is normal or not.

- The OPLEV dampings have been restored.

Attachment 1: TM_SUS.pdf  167 kB  | Hide | Hide all
TM_SUS.pdf
Attachment 2: TM_SUS_SD.pdf  67 kB  | Hide | Hide all
TM_SUS_SD.pdf
ELOG V3.1.3-